OpenStudy (anonymous):

Find d2y/dx2 in terms of x and y. y^4 = x^5

OpenStudy (anonymous):

y = \[y = \sqrt[4]{x^5} = x^{5/4}\]

OpenStudy (campbell_st):

well wouldn't you re-write the equation using index laws \[y = x^{\frac{5}{4}}\] so \[y' = \frac{5}{4} x^{\frac{1}{4}}\] then \[y" = \frac{1}{4} \times \frac{5}{4} x^{-\frac{3}{4}}\]